What is the age of Sipho and Litha in 6 years

1
Expert's answer
2022-03-07T17:26:06-0500

Let Sipho and Litha's present ages be x years and y years respectively.


We have that


"x+y=29" -----------> (1)


"x+6=y-3"


"=>x-y=-9" -----------> (2)


Adding equation (1) and (2) together, we have;


"2x=20"


"=> x = 10"

Put x=10 in equation (2)


"10-y=3"


"=> y=10-3=7"


Therefore, Sipho and Litha's present ages are 10years and 7years respectively.